A Food Marketing Institute found that 26% of households spend more than $125 a week on groceries. Assume the population proportion is 0.26 and a simple random sample of 324 households is selected from the population. What is the probability that the sample proportion of households spending more than $125 a week is less than 0.27?

Answer =  (Enter your answer as a number accurate to 4 decimal places.)
Answer #1

P - 26% = 0.26 n = 324, = 0.27 ² - P & P11-P)/n --- Pol & L0.27) = P. Pr (el = P(Z < 0-27–0-26 110-26x0-74)/324 Po (n L ) <0.